# NonGEMM Bench: Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads

### v3

@@ 3.2.2 End-to-End Inference Flow
The end-to-end inference flow profiles NonGEMM Bench models running inferences on data samples from the selected dataset, which consists of two stages. The first stage extracts the operator graph of a given model and preprocess input data. The second stage records the latency and energy of the inference runs.
Graph Extractor functions convert the model to be evaluated into a computational graph exposing the GEMM and non-GEMM operators for profiling. Based on the selected deployment flow, the graph extractor calls the corresponding APIs to trace through the model and generate the operator graph. For the ONNX Runtime deployment flow, the NonGEMM Bench harness calls ONNX graph exporters from the HuggingFace transformers library [44] or from the PyTorch libraries. For the PyTorch deployment flow, the graph extractor relies on the torch.fx [37] Tracer to obtain the computational graph. The Tracer propagates a proxy data structure into the model and captures the operators as graph nodes.
@@
Data Preprocessing functions are model specific functions that feed the profiling functions with input tensors. First, they fetch the raw data from the specified dataset and perform data cleaning. Empty sequences are removed from text data and images are re-scaled to each model’s specific input resolution. Second, they apply tokenization or image to tensor transformation to generate the input tensors.
End-to-End Inference Profiling Functions receive operator graphs and input tensors from the harness functions and run profiling iterations. The latency profiling functions record the operator breakdown of the end to end latency as well as the shape of the operators input tensors. Similar to the graph extractors, the latency profiling functions are deployment flow specific. For ONNX Runtime, we invoke Execution Provider profiling in the inference session. Currently, we only support operator level profiling for CPU and CUDA execution providers because profiling support to other ONNX RUNTIME backends is not stable yet. In PyTorch eager mode, we modify the torch.fx Interpreter to invoke the PyTorch profiler at each node of the operator graph. In TorchScript and TorchDynamo, we currently rely on the PyTorch Profiler. The energy profiling functions record the end to end inference power and energy consumption. We utilize system specific profiling tools to obtain CPU and GPU power profiles: NVIDIA’s System Management Interface (nvidia-smi) and AMD micro-Prof ($\mu$Prof) software profiling analysis tool.
3.2.3 MicroBench Flow
The MicroBench flow focuses on individual non-GEMM operators, which allows users perform deep analysis of their systems targeting non-GEMM operators. The flow consists of MicroBench Extractor, the NonGEMM Bench Operators Registry, and the MicroBench Operator Profiling (shown as ”u-Bench op Profiling in Figure 4).
MicroBench Extractor. This component populates the non-GEMM operators captured during end-to-end flow of NonGEMM Bench to the microbench operator registry. In addition to the operators, the registry stores the shape of input tensors to the non-GEMM operators as well as the implementation in their parent model. Such a step is crucial for keeping the microbenchmark realistic.
MicroBench Operator Profiling function (”u-Bench op Profiling” in Figure 4) fetch the non-GEMM operators from the registry and profile their performance utilizing latency/energy profiling in the end-to-end flow. For each operator, it generates input tensors using the recorded shapes and proceeds to evaluate the latency and the energy of the operator.
3.2.4 NonGEMM Bench Outputs
@@
NonGEMM Bench generates many statistics orgainzed into three categories: performance/cost, workload, and non-GEMM-specific reports. The performance/cost report includes key performance metrics such as the end-to-end latency with operator level break-downs, energy, and peak memory usage. The workload report includes the shape of the tensors captured during inference on realistic data. Finally, the non-GEMM report provides insights on non-GEMM operators such as the number of operator variants of the same class of non-GEMM operator (e.g., layernorm implementation could have multiple variants for each ML framework), non-GEMM operator trace on different domains, and so on.
@@
Table 2: Non-GEMM operators in eight popular model variants selected from Hugging Face listed in Table 1.
@@
Arithmetic Add Segformer [2, 16384, 32]
@@
4 Case Studies
We perform a thorough analysis of the performance data generated by running NonGEMM Bench on multiple hardware platforms show in  Table 3. We collect the data by running NonGEMM Bench in PyTorch eager mode on three systems configuration ranging from mobile CPUs to data center grade CPUs and GPUs to capture the behavior of the operators when subjected to various acceleration capabilities. We notice that for most of the models, the time spent to compute non-GEMM operators increases noticeably with respect to the total execution time when running the model on the GPU.
Table 3: Hardware platform configurations used for NonGEMM Bench case studies.
Class CPU GPU
Data Center AMD EPYC 7763 NVIDIA A100
Workstation Intel i9-13900K NVIDIA RTX 4090
Mobile Intel i7-13700HU NVIDIA RTX 4060m
4.1 Impact of Non-GEMM on Each Task Domain
We discuss how non-GEMM operator’s contribution to the end-to-end inference and energy changes depending on task domains.
4.1.1 Image Classification
As shown in Figure 5, we observe a clear change in relative break-down latency across operators when using a GPU. For example, the normalization operation is minor before GPU acceleration, but it becomes the dominant operator with a GPU, accounts for over 50% of the total execution time, on average. We observe similar patterns for all the other cases: GEMM operators is dominant on CPU-only systems, but non-GEMM operators become significant or considerable after GPU acceleration. For the batch size of 1, the relative execution time of Non-GEMM operators for running ViT-b16 and ViT-l16 increases from 23.5% and 18% when running on a CPU only platform to 60% and 55%, respectively, on the GPU as shown in  Figure 5. Similarly, for the workstation GPU system, the non-GEMM operators execution time proportion increases to 56% and 45% for ViT-b16 and ViT-l16 to almost split the total execution time evenly with GEMM operators. For the three Swin transformers with a batch size of 1, the non-GEMM operators contribution to the overall latency increases as well when running the models on the GPU. The small Swin transformer model spends 55% of the computation time executing non-GEMM operators on the data center GPU system compared to a modest 33% contribution on the CPU only system. On the workstation GPU system, the small Swin transformer also computes non-GEMM operators for 57% of the time, a 2x increase compared to the workstation CPU system.
Impact of the batch size. When increasing the batch size, the latency distribution across GEMM and non-GEMM operators of IC models running on GPU follows the same trend as that of single batch. However, this trend is less noticeable for the largest vision transformer, ViT huge (632M parameters). For a batch size of 8, the non-GEMM operator latency is 43% and 51% for ViT-l16 and base Swin transformer respectively (Figure 5) running on the data center GPU, an 3.9$\times$ and 3$\times$ increase compared to the CPU only results. However, the huge vision transformer only shows a 1.25x increase in the latency proportion of non-GEMM operators when accelerating with a GPU.
The non-GEMM operators with the most significance to the execution time of classification models are presented in Figure 5. As shown in  Table 4, Normalization operators are the most expensive non-GEMM operator in classification models accounting for an average of 17% of the execution time on the data center GPU, followed by Arithmetic operators with 14% (Figure 5). Based on NonGEMM Bench collected data, image classification models require  48% of execution time on non-GEMM operators on GPU systems, on average, across all the evaluated models.
4.1.2 Object Detection
Non-GEMM operator’s latency is also noticeable for object detection with a GPU. On the data center, non-GEMM operators dominated the execution time of FasterRCNN by consuming 82% of the total inference time when we accelerate the workload compared to 46% when running the CPU only (Figure 5). DETR and MaskRCNN showed a similar behavior after acceleration, with a 2x increase in the time spend on non-GEMM operators. DETR computed non-GEMM operators for 77% of the inference time on the data center GPU, while MaskRCNN spent 80% on the data center and workstation GPUs. For object detection models, Normalization operators exhibit the highest execution time followed by activation and memory operators as indicated in (Figure 5). All detection models in NonGEMM Bench use normalization functions with custom implementations that do not leverage available library normalization kernels. This leads to additional overhead to execute the normalization function. Overall, the evaluation data shows that detection models get dominated by non-GEMM operators (78%) when executing the workloads on a GPU system.
4.1.3 Segmentation
For segmentation models, the non-GEMM operators account for 47% of the execution time when running the inference on the CPU on average. Upon running the model on the GPU, this number slightly increases to 53%. Figure 5 shows memory and arithmetic operators to be the most latency expensive non-GEMM operator on GPU accounting for 17% and 12.5% respectively. For the SegFormer model, Normalization operators are the most expensive (15% of the execution time), while Memory operators dominate the non-GEMM operators of Maskformer (24%).
4.1.4 Language Models
GPT2 and Bert executed non-GEMM operators for 77% and 57% of the time when running on a single batch as indicated in Figure 5. For GPT2-large, the non-GEMM operators clearly dominated the execution time on GPU systems as well. Across the three GPU HW classes, non-GEMM operators occupy 76% of execution time on average. Table 4 show that Activation (23% on GPT2) and element wise arithmetic (20% on Bert) operators are the most expensive non-GEMM operators in language models running on GPU.
Table 4: NonGEMM Bench Most Expensive NonGEMM Operator Group for selected NonGEMM Bench Models Running on Data Center GPU Platform.
Models Model Alias Batch Size Operator Group % of Exec. Time
ViT Base 16 Vt-b 1 Norm 20
8 Norm 17
ViT Large 16 Vt-l 1 Norm 20.5
8 Arith 17
Swin Tiny Sw-t 1 Norm 19
8 Norm 18.5
Swin Small Sw-s 1 Norm 18
8 Norm 17
Swin Base Sw-b 1 Norm 18
8 Norm 18
FasterRCNN FRCNN 1 Norm 60.5
2 Norm 59
8 Norm 49
MaskRCNN MRCNN 1 Norm 58
2 Norm 56
8 Norm 49
DETR DETR 2 Norm 40
GPT2 gpt2 1 Act 23
64 Act 16
GPT2 X-Large GPT2-XL 1 Act 23
64 Act 9
Llama2-7B Llama2 1 Arith 23
Bert bert 1 Arith 20
64 Arith 10
@@
Figure 6: Breaking down the execution time of models running on the Workstation (CPU only) and (CPU + GPU) configurations across operator groups.
Large language model’s latency is dominated by GEMMs when running on Data center, consisting of 79% of the total inference time on average across platforms. When utilizing the GPU, the GEMM latency no longer bottlenecks the total execution time. The GEMM operators time drops to 30% and 22% for Llama2-7b and GPT2-XLarge respectively with non-GEMM operators executing for 74% on the GPU configuration. Figure 7 (a) suggests that the bottleneck are activation, arithmetic, and memory operators. Element-wise arithmetic operators are the most time-consuming non-GEMM operators for Llama2-7b, accounting for 23% of the total runtime (Table 4, while activation functions occupy 23% of GPT2-XL total execution time.
GPT2 models implement their own GELU activation in the Hugging Face library which does not have a direct mapping to a GELU GPU kernel when running in Pytorch eager mode. This leads to additional overheads due to running multiple kernels to compute the activation. Similarly, Llama uses a normalization function that does not map to a GPU kernel in eager PyTorch which leads to similar overheads that increase the latency of the non-GEMM operators.
4.2 Impact of Deployment Flow on Non-GEMM
Running in eager PyTorch, Non-GEMM operators occupy large proportions of the execution time for various applications running on the GPUs , as summarized in Table 4. Nevertheless, deployment frameworks (e.g. [16, 11]) are typically used to improve the performance beyond the eager mode [6]. They apply transforms to the computational graph level and leverage accelerators (if available) to optimize the inference runs. To understand how deployment frameworks affect inference performance distribution across operators, we evaluate NonGEMM Bench models with ONNX Runtime (ORT). ORT is a cross platform ML framework that accelerates ML model inferencing [16]. It leverages hardware accelerators (e.g. GPUs) and operator graph level optimizations to improve the performance of ML models. We run selected models with the CUDA execution provider and observe the following results.:
Non-GEMM operators account for a significantly larger proportion of the total execution latency (Figure 8) when running on different grade GPUs. This is aligned with the trend observed with the PyTorch eager execution on GPU. On average, the time spend on non-GEMM operators increases from 52% to a 73% after deploying the models on GPU with the ORT cuda execution provider.
Memory operators almost dominate the execution time in three application domains, classification, detection and language models by occupying 66%, 44%, 59 % respectively. For instance, memory operators are the most latency expensive non-GEMM operators in GPT2-XLarge and Llama2-7b as they consume 76% and 59% of the total inference time respectively (Figure 7). In detection models, memory operators consume 37% of the total inference time in ORT as well (Figure 8).
One reason leading memory operators to dominate the inference latency is their high frequency, notably in the large language models. They consist of 80% and 62% of the total operators in the GPT2-XLarge and Llama2-7b . Consequently, the cumulative latency to run these operators becomes the inference latency bottleneck. Another reason is the limitations in operator support for GPU deployment in ORT. Memory operators are a subset of non-GEMM operators that the cuda execution provider does not fully support, hence it assigns them to be computed on the CPU. In the case we hit an unsupported operator while executing on the GPU, the runtime falls back to CPU to carry the computation. As a result, additional overheads to transfer the data between devices are added to the operators execution time that increases the latency of the fallen-back operators, which aggravates frequently used operators.
Despite dominating the execution time in ORT, memory operators were not the main non-GEMM bottlenecks PyTorch eager mode for the evaluated models. Figure 7 highlights the differences in operator latency distribution for GPT2-XLarge and Llama2-7b running in PyTorch eager mode and ORT. This indicates that the deployment flows have distinct effects on the execution of non-GEMM operators in an end to end inference.
@@ 4.2 Impact of Deployment Flow on Non-GEMM
Figure 7: The impact of deployment software toolchain into the latency breakdown. (a) PyTorch (Eager Mode) and (2) ONNX Runtime on a data center class GPU (A100).
Figure 8: The latency breakdown of the inference using ORT on two platform configurations: mobile and Data Center.
4.3 Key Observations and Insights

### v4

@@ III-B2 End-to-End Inference Flow
The end-to-end inference flow profiles NonGEMM Bench models running inferences on data samples from the selected dataset.
First, the Graph Extractor extracts the operator graph of a given model based on the selected deployment flow. NonGEMM Bench calls graph exporters from the HuggingFace transformers library [62] or from the PyTorch libraries.
@@
Second, model specific Data Preprocessing functions fetch the raw data from the specified dataset, clean the data, and apply tokenization or image to tensor transformations to generate input tensors for the profiling stage.
Third, End-to-End Inference Profiling functions run the profiling iterations on the received the model and its inputs. The profiling functions are deployment flow specific. For ONNX Runtime, we invoke Execution Provider profiling in the inference session. Currently, we only support operator level profiling for CPU and CUDA execution providers [17] because profiling support to other ONNX RUNTIME backends is not stable yet. For PyTorch, we use the PyTorch Profiler [49]. For TensorRT, we leverage the TensorRT open source software framework profiling APIs [44].
III-B3 NonGEMM Bench Outputs
@@
NonGEMM Bench generates many statistics organized into three categories: performance, workload, and non-GEMM-specific reports. The performance report includes key performance metrics such as the end-to-end latency with operator level break-downs and peak memory usage. The workload report includes the types of operators and the shape of the tensors for each operator captured during inference on realistic data. Finally, the non-GEMM report provides insights on non-GEMM operators, such as the number of operator variants of the same class of non-GEMM operator (e.g., layernorm implementation could have multiple variants for each ML framework) and non-GEMM operator trace on different domains.
@@
TABLE II: Non-GEMM operators in eight selected model variants from Table I and their characteristics.
@@
Elmt-wise Arithmetic Add Segformer [2, 16384, 32]
@@
IV Case Studies
We perform a thorough performance analysis of models in Table I on multiple hardware platforms listed in  Table III. The evaluated hardware platforms span workstation and data center-class CPUs and GPUs. We employ PyTorch, ONNX runtime, and TensorRT as deployment flows to investigate their impact on the performance horizon.
Using the aforementioned workload, hardware platforms, and deployment flow, we characterize the GEMM/non-GEMM performance for each model in NonGEMM Bench and investigate the impact of the hardware platform and deployment flow. In addition, we also investigate the impact of widely adopted optimizations (operator fusion and quantization) and show that the non-GEMM performance is an issue even after applying common optimizations.
TABLE III: Hardware platform configurations used for case studies.
ID Class CPU GPU
Device Device Mem. TOPS
A Datacenter AMD EPYC 7763 Nvidia A100 80 GB 624
B Workstation Intel i9-13900K Nvidia RTX 4090 24 GB 660
IV-A Non-GEMM Performance Characterization Results
We first conduct a performance characterization study using PyTorch and present the characterization results on datacenter and workstation platforms (Platforms A and B in Table III) in Figure 5. We focus on those two platforms since some models (e.g, large language models) do not fit into the mobile devices. As expected, the relative contribution of non-GEMM operators to the end-to-end latency significantly increases after GPU acceleration, on average, from 17.2% to 42.3%. However, we observe each model show different trends, mainly affected by the non-GEMM operator types and the number of GEMM and non-GEMM operators in the original model. We summarize the most time-consuming non-GEMM operators in Table IV from the datacenter class platform (Platform A), which shows the diversity of the dominating non-GEMM operators in each model. We highlight some notable models in each task category and delve into the details of their non-GEMM performance.
IC Task: Swin Transformer. For every Swin Transformer variant(Sw-t, -s, and -b), the memory operator group is the most latency-dominant non-GEMM operator group, which accounts for 32% of the total latency, on average, on data center platform with GPU acceleration. Those memory operators originate from the Swin Transformer’s unique window shape (cross-shaped) [37], which is not well-aligned with memory layout of tensor data organized in dimension orders.
OD Task: DETR. After GPU acceleration, DETR has shown significant non-GEMM operator presence, which accounts for 65.8% of the total latency, on average. The major source of the non-GEMM latency is in the normalization operators, whose percentages are 35% and 32% on Platforms A and B, respectively. We observe the normalization functions are based on a custom implementation, which are identified as independent kernel. Kernel launch overheads accumulated for independent runs for the custom normalization function significantly contribute to the non-GEMM latency. However, we also observe that an advanced deployment software (TensorRT) can fuse those operators and significantly improve the non-GEMM performance. We discuss the details later in Section IV-B.
IS Task: Maskformer. MaskFormer utilizes Swin Transformer as its backbone, which introduces many memory operators as we discussed for Swin Transformer. As a result, memory operator is the dominant non-GEMM operator, which accounts for 40.8% of the total latency, on average, as we can observe in Figure 5 (f) and (n).
NLP Task: GPT2. The latencies of non-GEMM operators in GPT2 variants are considerable, which account for 45.0%, on average, as we observe in Figure 5 (h) and (p). The dominant non-GEMM operator is activation, GELU, which required 26.4% of the total latency.
TABLE IV: Most time-consuming non-GEMM operator groups for selected Models (Platform A, with GPU acceleration, average across batch sizes).
Task Domain Model Operator Group Latency Percentage (%)
Image Classification Vt-b Norm 14.0
Vt-l Norm 13.3
Vt-h Norm 11.2
Sw-t Memory 31.8
Sw-s Memory 33.1
Sw-b Memory 32.8
Object Detection FRCNN Elmt-wise Arith. 34.4
MRCNN Elmt-wise Arith. 33.6
DETR Norm 34.8
Image Segmenation MF Memory 40.8
Seg Normalization 17.4
NLP gpt2 Act 30.2
GPT2-L Act 29.9
GPT2-XL Act 28.1
Llama2 Norm 14.9
bert Norm 13.1
Summary. We observe GPU acceleration significantly increase the percentage of non-GEMM operators and increase the importance of non-GEMM operators in performance optimization. Also, the dominant operators are diverse in each model, which indicates an optimization technique tailored for a single operator cannot fully address the non-GEMM performance problem.
@@ IV-A Non-GEMM Performance Characterization Results
IV-B The Impact of Deployment Flow on Non-GEMM Performance
Deployment flows such as ONNX runtime [17] and TensorRT [43] are widely used for serving model inferences. Such flows apply various optimizations to each model, which includes the computational graph level optimizations (e.g., operator fusion) and backend assignment (e.g., utilizing Tensor Core in Nvidia GPUs). To understand the effect of deployment frameworks on the operator performance, we conduct two case studies (1) comparing PyTorch and ONNX Runtime (ORT) results on LLMs(focus: general optimizations w/o operator fusion) on LLMs, and (2) comparing PyTorch and TensorRT results (focus: operator fusion).
IV-B1 Case Study 1: Non-GEMM Performance on LLMs across PyTorch and ORT
We run two LLMs (GPT2 and Llama2) using the CUDA execution provider in ORT and compare their non-GEMM performance percentage in Figure 6.
In the results, we observe the presence of non-GEMM operators significantly increase, from 52.6% to a 80.75%, on average. We observe the percentage of memory operators significantly amplifies after applying ORT, from 3.2% to 66.8%. Such a result originates from the limitation in operator support in ORT. Many memory operators in the evaluated LLMs are not supported by the cuda execution provider in ORT, which leads to inefficient execution on CPUs involving data transfer between a CPU and a GPU. Combined with the high frequency of those operators, the inference latency of memory operators increases significantly, as shown in Figure 6 (b). The results imply two insights: (1) Model deployment flow can significantly aggravate the non-GEMM performance challenge and (2) the dominant non-GEMM operators differ depending on the operator support in a deployment flow.
@@ IV-B1 Case Study 1: Non-GEMM Performance on LLMs across PyTorch and ORT
Figure 6: The impact of deployment software toolchain into the latency breakdown on language models. (a) PyTorch and (2) ONNX Runtime on a data center class GPU (A100).
Figure 7: The latency of non-GEMM operator in (a) Swin-b, (b) DETR, and (c) Segformer is still considerable despite applying operator fusion in TensorRT. The numbers at the top of each bar represent the latencies in milliseconds.
IV-B2 Case Study 2: Non-GEMM Performance with Operator Fusion
Operator fusion is a key optimization technique for accelerating inference workloads [14, 25, 57, 42, 43, 30]. Fusion combines multiple operators in a single kernel to reduce the number of costly kernel launches and minimize the number of redundant offchip memory accesses [42]. TensorRT [43] is a widely adopted inference framework released by Nvidia that applies the operator fusion technique targeting GPUs. Operator fusion in TensorRT detects specific patterns (e.g., three consecutive element-wise operators [43]) in the operator graph and fuses nodes captured in the patterns to enhance inference performance by reducing redundant memory accesses around non-GEMM operators.
Impact of Operator Fusion. To understand the impact of operator fusion on non-GEMM performance horizon, we conduct a case study on four models, which compares the TensorRT (with fusion) and PyTorch (without fusion). We present the results in Figure 7, which shows the inference latency breakdown between GEMM and non-GEMM operators. We notice that fusion does not completely address the non-GEMM bottleneck in Swin-b model. For example, the percentage of non-GEMM operators in the total latency changes from 56.2% to 32.2%, on average, after operator fusion by TensorRT. The reduction in the percentage is based on the improvement of non-GEMM performance via operator fusion, which reduces 88.6% of latency after fusion, as summarized in Table V. However, the non-GEMM operators still account for 32.2% of total latency, which is considerable. This shows that operator fusion cannot eliminate the non-GEMM performance challenge and motivates further studies toward non-GEMM performance.
Factors Affecting the Effectiveness of Fusion. Although most results indicate considerable impact of non-GEMM even after operator fusion, we observe operator fusion on DETR model is exceptionally effective. Therefore, we conduct a deep-dive study, investigating the percentage of fused non-GEMM operators (i.e., fusion rate) and performance improvements after fusion, as listed in Table V. We observe the strong non-GEMM performance improvements for DETR originates from high fusion rate of 30%, which led to 13.5$\times$ non-GEMM speedup. This leads to the large percentage reduction of non-GEMM in the total latency, from 67% to 18.5%, on average.
However, the fusion rate is not the only factor that determines the non-GEMM speedup. For example, DETR and Segformer have similar fusion rates (30% andn 27%, respectively), but the amount of non-GEMM performance improvements are significantly different: 13.5$\times$ and 2.39$\times$, respectively. We analyze the execution trace and find that batch normalization is the main source of the differences. The main difference is the fusion pattern: most batch normalization operators (100% of total) in DETR were fused with GEMM-operators (CONV+BN+ReLu pattern) while those in Segformer were fused with other non-GEMM operators (97.8% of total). The results indicate that the effectiveness of operator fusion relies on not only the overall fusion rate but also the fusion patterns. Our observation confirms that the operator fusion cannot fully address the non-GEMM performance challenge, even if it can be very effective on some patterns.
IV-C Non-GEMM Performance with Quantization
Quantization refers to the model optimization technique, which reduces the bit precision of model weights and/or activations to enhance the computational performance and efficiency of DNN inference. Quantization is a widely-adopted technique [28, 40] including heavy models like LLMs [19, 35, 63, 16]. For example, LLM.int8() [16] introduced an 8-bit - 16-bit mixed precision quantization scheme for LLMs. LLM.int8() is a state-of-the-art quantization method, which quantizes more than 99% of the linear layers in OPT LLM to an 8-bit precision. Therefore, we adopt LLM.int8() and characterize GEMM and non-GEMM performance to understand the impact of quantization on the non-GEMM operator performance problem.
We conduct a case study on quantization using LLM.int8() on Llama3 LLM on the datacenter platform with GPU acceleration and present the results in Figure 8. In the results, we observe that non-GEMM operators dominate the latency after quantization, changing from 29.3% to 76.7%, on average. Such a significant shift in the latency distribution is mainly based on the GEMM performance improvements from 8-bit arithmetic, which reduced the latency by 38.2%, on average. However, based on our analysis on the execution traces, non-GEMM performance aggravates because the 8-bit data need to be dequantized and re-quantized for non-GEMM operators, which requires 16-bit floating point arithmetic. This led to 6510 additional non-GEMM operators into the computational graph, which led to a significant increase of the non-GEMM latency by 5.6$\times$ after quantization. Combined together, the overall percentage of non-GEMM operators in the total latency dominate after quantization, which makes non-GEMM operators as the major optimization target.
In the case study on Llama3 8B, we observe the element-wise arithmetic operators originate from dequantization/requantization (DQRQ) process dominate in the inference latency, which adds 20% extra non-GEMM operators to the original computational graph. Also, we observe longer sequence length leads to higher percentage in the element-wise arithmetic operators. For example, as we increase the sequence length from 512 to 8192, the latency percentage of element-wise arithmetic operators increase from 31.8% to 63.8%. As current trends in the LLMs are toward longer sequence lengths [18, 5], the non-GEMM performance issues in longer sequences originating from DQRQ costs will aggravate, which motivates efforts in non-GEMM performance optimization.
Figure 8: Non-GEMM operators latency contribution to the inference of a quantized Llama3 8B using 8 bit quantization [16]. The numbers at the top of each bar represent the latencies in milliseconds.
TABLE V: The non-GEMM latency before and after applying fusion with TensorRT. The values between brackets represent the percentage w.r.t the total inference latency.
Model Non-GEMM Non-GEMM Latency
Fusion Rate Before Fusion After Fusion
Swin-t 8.75% 7.53 ms (56.4%) 0.97 ms (39%)
Swin-b 7% 14.59 ms (58%) 1.65 ms (32.25%)
DETR 30% 32.17 ms (66.4%) 2.38 ms (18.5%)
Segformer 27% 5.57 ms (41%) 2.33 ms (41%)
IV-D Key Observations and Insights